OpenStudy (anonymous):

Given a point and the slope, write an equation of the line in point-slope form. m = -4/5 (-4, 9) i dont understand this.

OpenStudy (anonymous):

y-9=-4/5(x+4)

OpenStudy (anonymous):

im sorry is there any fuller explaination?

OpenStudy (anonymous):

the formula is y-y1=m(x-x1) where x1, y1 r ur coords and m is ur slope

OpenStudy (anonymous):

Oh Thank you!!

OpenStudy (anonymous):

yup.